OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of three i (i already found the answer i just need someone to verify if it is right)

OpenStudy (jdoe0001):

\(\bf -3+3\sqrt{3}\ i\quad ?\)

OpenStudy (jdoe0001):

\(\bf \begin{array}{ccllll} -3&+3\sqrt{3}\ i\\ \uparrow &\uparrow \\ {\color{brown}{ a}}&{\color{brown}{ b}}\\ \end{array}\qquad \begin{array}{llll} r=\sqrt{{\color{brown}{ a}}^2+{\color{brown}{ b}}^2}\\ r=\sqrt{(-3)^2+(3\sqrt{3})^2}\\ r=\sqrt{9+3^2\cdot 3}\\ r=\sqrt{36}\to 6 \end{array} \\ \quad \\ \theta=tan^{-1}\left(\cfrac{\cancel{ 3 }\sqrt{3}}{-\cancel{ 3 }}\right)\implies \theta=tan^{-1}(-\sqrt{3}) \\ \quad \\ thus\implies r[cos(\theta)+i\ sin(\theta)]\)

OpenStudy (anonymous):

6(cos(2pi/3) + i sin(2pi/3))

OpenStudy (anonymous):

is that right?^

OpenStudy (jdoe0001):

hmmm well for \(\bf tan^{-1}(-\sqrt{3})\) I get -60 degrees or \(\bf -\cfrac{\pi}{3}\to \cfrac{5\pi}{3}\)